# Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings

### v2

@@ 3.1 Dataset Curation
Video annotation. We cross-reference the video titles with the predefined list of procedures. When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as specified in Sec. B of the supplementary material. All annotations are manually reviewed and validated to ensure the accuracy of the labels assigned.
4 Proposed SurgFM Foundation Model
In this section, we describe how we build SurgFM, our foundation model trained on Surg-3M, and motivate its design choices. This section includes two ablation studies to identify the optimal self-distillation configuration and the most suitable backbone architecture for surgical videos; a description of our image preprocessing pipeline, and our proposed augmented distillation mechanism with its corresponding ablation study. The results of the ablation experiments are presented in Table 2.
Self-supervised methods. We examine the impact of the different self-supervision mechanisms (Table 2). We evaluate DINO [8], DINOv2 [46], and MAE [23]. For these experiments, we trained the models in their standard configurations, a ViT-B backbone [14] pretrained on ImageNet-1k [13] with frozen weights and a linear classification head. Stemming from these results, which align with those in the literature [49], DINO consistently outperforms DINOv2 and MAE on all downstream tasks, with surgical tool detection and action recognition showing a large margin. Therefore, we adopt DINO’s self-distillation [8] as our pretext task for self-supervision.
Backbone models. As noted in [56], ViTs [14] face challenges to capture fine details, which is essential to solving tasks such as tool presence detection and surgical action recognition. Given this, we opted to evaluate the performance of ViT-L [14] against a modern convolutional architecture, ConvNeXt-L [36]. Both were pretrained on Surg-3M and fine-tuned on Cholec80 [54] and CholecT50 [42], for surgical phase recognition, tool presence detection, and surgical action recognition. The results are shown in Table 2. ConvNeXt-L outperformed ViT-L, suggesting that convolutional networks are indeed better equipped to capture those fine image details that are relevant to solve surgical tasks. Therefore, we use ConvNeXt-L as the backbone of our SurgFM model.
Image preprocessing pipeline. Surgical videos often exhibit superimposed manufacturer-specific user interface (UI) information generated by the software of the surgical stack. While certain researchers might derive value from analyzing these UI elements (e.g., some interfaces provide contextual information regarding the surgical instruments utilized in the video), a foundational model for surgery should prioritize learning correlations between tissue appearance and tool-tissue interactions. This focus is essential to avoid acquiring spurious correlations with manufacturer-specific UI elements. Therefore, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 Nano model [31] to detect and crop out UI content (see Fig. S2 in supplementary material).
[…]
Method Type Cholec80 AutoLaparo M2CAI16 GraSP
Accuracy (%) Precision (%) Recall (%) Jaccard (%) F1-score (%) Accuracy (%) Precision (%) Recall (%) Jaccard (%) Accuracy (%) Precision (%) Recall (%) Jaccard (%) mAP (%)
ConvNeXt-L [36] G $91.4\pm 6.7$ $92.3\pm 3.8$ $90.5\pm 7.7$ $82.1\pm 7.1$ $88.6\pm 6.8$ $82.3\pm 9.7$ $78.3\pm 21.9$ $69.9\pm 31.9$ $60.5\pm 27.2$ $88.3\pm 10.0$ $89.4\pm 4.7$ $89.5\pm 3.3$ $78.4\pm 6.2$ $73.4$
EndoNet [54] S $81.7\pm 4.2$ $73.7\pm 16.1$ $79.6\pm 7.9$ $76.5$
MTRCNet-CL [27] $89.2\pm 7.6$ $86.9\pm 4.3$ $88.0\pm 6.9$ $87.4$
PhaseNet [52] $78.8\pm 4.7$ $71.3\pm 15.6$ $76.6\pm 16.6$ $73.8$ $79.5\pm 12.1$ $64.1\pm 10.3$
SV-RCNet [26] $85.3\pm 7.3$ $80.7\pm 7.0$ $83.5\pm 7.5$ $82.1$ 75.6 $64.0$ $59.7$ $47.2$ 81.7 $\pm$ 8.1 81.0 $\pm$ 8.3 81.6 $\pm$ 7.2 65.4 $\pm$ 8.9 -
OHFM [62] $87.3\pm 5.7$ $67.0\pm 13.3$ $85.2\pm 7.5$ $68.8\pm 10.5$
TeCNO [12] $88.6\pm 7.8$ $86.5\pm 7.0$ $87.6\pm 6.7$ $75.1\pm 6.9$ $77.3$ $66.9$ $64.6$ $50.7$ $86.1\pm 10.0$ $85.7\pm 7.7$ $88.9\pm 4.5$ $74.4\pm 7.2$
TMRNet [29] $90.1\pm 7.6$ $90.3\pm 3.3$ $89.5\pm 5.0$ $79.1\pm 5.7$ $78.2$ $66.0$ $61.5$ $49.6$
AVT [20] - - - - - $77.8\pm 9.4$ $68.0$ $62.2$ $50.7$
Trans-SVNet [30] $90.3\pm 7.1$ $90.7\pm 5.0$ $88.8\pm 7.4$ $79.3\pm 6.6$ $78.3$ $64.2$ $62.1$ $50.7$ $87.2\pm 9.3$ $88.0\pm 6.7$ $87.5\pm 5.5$ $74.7\pm 7.7$
LoViT [35] $92.4\pm 6.3$ $89.9\pm 6.1$ $90.6\pm 4.4$ $81.2\pm 9.1$ $81.4\pm 7.6$ $85.1$ $65.9$ $55.9$
SlowFast± [17] - - - - - - - - - - - - - $70.1$
TAPIR± [55] - - - - - - - - - - - - - $74.6$
TAPIS± [4] - - - - - - - - - - - - - $\mathbf{76.1}$
Ramesh et al. [49] F - - - - $81.6$
SSL4EVA [24] - - - - $86.9$
EndoViT [5] 89.4 - - - - - - - - - - - - -
SurgFM (ours) $87.0\pm 10.3$ $87.3\pm 7.5$ $87.7\pm 4.3$ $75.7\pm 9.2$ $83.1\pm 11.0$ $81.7\pm 10.9$ $\mathbf{86.3\pm 8.9}$ $69.8\pm 30.9$ $60.2\pm 26.1$ $82.6\pm 10.7$ $85.5\pm 9.9$ $87.3\pm 5.9$ $73.2\pm 9.4$ $66.8$
SurgFM (ours) $\mathbf{92.7\pm 8.3}$ $\mathbf{93.0\pm 4.5}$ $\mathbf{93.1\pm 4.7}$ $\mathbf{85.1\pm 7.9}$ $\mathbf{90.6\pm 6.7}$ $\mathbf{85.2\pm 9.9}$ $82.9\pm 17.3$ $\mathbf{73.2\pm 32.0}$ $\mathbf{64.8\pm 27.3}$ $\mathbf{89.9\pm 7.1}$ $\mathbf{89.6\pm 6.5}$ $\mathbf{89.9\pm 6.9}$ $\mathbf{79.4\pm 9.9}$ $75.9$
Proposed augmented distillation method. Inspired by DINO [8], we train a student network $f_{\boldsymbol{\theta_{s}}}$ to solve for the mapping $\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))=\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ where $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ is an input image, $\sigma$ is the softmax function, $f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})\in\mathbb{R}^{C}$ is a target vector of dimension $C=2^{16}$ estimated from a teacher network $f_{\boldsymbol{\theta_{t}}}$, and $\boldsymbol{\theta_{s}}$ and $\boldsymbol{\theta_{t}}$ are the parameters of the student and teacher models, respectively. We use the outputs of both networks to build two categorical distributions, $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ and $P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))$, where $\sigma_{z}$ denotes the $z^{th}$ element of the softmax output, and $z\in\{1,\dots,C\}$. In training, we minimize the negative log-likelihood of $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))$ following
$\small\min_{\boldsymbol{\theta_{s}}}\sum_{i=1}^{I}$ $\sum_{\begin{subarray}{c}\boldsymbol{u}\in U_{i}\\ \boldsymbol{v}\in V_{i}\cup W_{i}\\ \boldsymbol{u}\neq\boldsymbol{v}\end{subarray}}\Big{[}d(\boldsymbol{u},\boldsymbol{v})-n(\boldsymbol{u},\boldsymbol{v})\Big{]}$ (1)
where $d(\boldsymbol{u},\boldsymbol{v})=\log\sum_{k^{\prime}=0}^{C-1}\exp\left(\sigma_{k^{\prime}}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))\right)$, $n(\boldsymbol{u},\boldsymbol{v})=\sigma_{\arg\max_{z}P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{u})))}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))$, $I$ is the number of images in the dataset, $U_{i}$ contains two global augmentations of $\boldsymbol{x_{i}}$, and $V_{i}$ contains two global augmentations and four augmented crops of $\boldsymbol{x_{i}}$. A foundation model for surgery should learn invariance to 1) minor movements of instruments and tissues observed between adjacent video frames, and 2) minimal appearance changes across patients undergoing the same surgical procedure (e.g., organ color variations). Diverging from DINO  [8], and relying on the annotations available in Surg-3M, we introduce $W_{i}$ to provide a supervisory signal that encourages the model to learn these invariances. $W_{i}$ contains two images, and four augmented crops of these two images. The procedure for selecting these two images is depicted in Fig. 4. We use the $k$-nearest neighbors (in embedded space) of the input image from other videos of the same procedure type to learn invariance to appearance changes across patients, and the prior and subsequent video frames of the input image to learn invariance to the motion occurring between consecutive frames.
We evaluate the performance of models trained with and without our augmented distillation (see Table 2) on surgical phase recognition (Cholec80), tool presence detection (Cholec80), and surgical action recognition (CholecT50). With augmented distillation, F1-score improved by $>1pp$ in both tool detection and action recognition tasks.
Following these ablation results, we incorporated augmented distillation in our final SurgFM model.
@@ 4 Proposed SurgFM Foundation Model
Figure 4: Proposed augmented distillation method. To encourage invariance in SurgFM to minor surgical motion and subtle appearance changes across surgeries, we introduce $W_{i}$ (Eq. 1). The key component of $W_{i}$ is a pair of images (c) that are randomly selected from an augmentation pool (b). To populate the augmentation pool, which has capacity for four images, we first retrieve the nearest neighbors(1) of the input image from other videos of the same procedure type, but only include them in the pool if the cosine distance to the input image is smaller than 3×3$\times$3 × the distance between the input image and its preceding frame in the video that it was extracted from. When not enough suitable neighbors are found, we supplement the pool with adjacent video frames(2)(3).
@@
Proposed video classification model. Considering that each surgical procedure is localized to a specific region of the body, we posit that a characteristic scene can be associated with each procedure. Based on this assumption, and inspired by active learning [21], we propose SurgFM-Vid, a video classification method that employs SurgFM embeddings for video classification by aggregating them following
$\small\boldsymbol{v_{e}}=\sum_{j=0}^{J-1}\omega_{j}\boldsymbol{\phi}_{j}$ (2)
@@
5 Evaluation
5.1 Downstream Tasks, Datasets and Metrics
@@
We evaluate SurgFM on three surgical downstream tasks: surgical phase recognition [54, 53, 57, 4], surgical tool presence detection [54, 4], and surgical action recognition [48, 4]. To do so, we use the most widely adopted datasets, data splits, and evaluation metrics.
Surgical phase recognition. We use M2CAI16 [54], Cholec80 [54], AutoLaparo [57], and GraSP [4]. For this task, we report accuracy, precision, recall, Jaccard, F1-score, and mean Average Precision (mAP). The accuracy for a video is computed by dividing the number of frames whose class has been correctly predicted by the total number of frames. The overall accuracy for a dataset is defined as the mean accuracy value over all videos, henceforth referred to as video-level accuracy [28, 29, 35, 27]. Precision, recall, and Jaccard are computed differently than accuracy [28, 29, 35, 27]. These three phase-level metrics are computed for each class and video, then averaged: first within videos, and secondly across classes. The F1-score, computed for Cholec80, is also computed for each class and video, but averaged first over classes and then over videos, and referred to as video-level F1-score [49, 24]. For GraSP, the frame-level mAP was calculated by averaging the mAP scores across all frames in the test set [4].
Surgical tool presence detection. Cholec80 and GraSP are utilized in the surgical phase recognition task. We report mAP across all frames in the testing dataset [49, 4].
Surgical action recognition. For this task we use CholecT50 [44] and SAR-RARP50 [48]. For CholecT50, we report the video-level mAP, which is computed by first calculating the average mAP for each class across all videos and then taking the overall mean of these class averages [44]. For SAR-RARP50 [48], we report video-level accuracy and video-level F1@10 [48].
5.2 Downstream Task Evaluation Protocols
To capture temporal correlations in the surgical phase recognition downstream task, we couple our SurgFM backbone with a TCN head [12]. This TCN head takes both current and previous frames into account, but excludes future frames. For a fair comparison with previous methods [26, 54, 30, 12, 27, 35], we run our evaluation in Cholec80 and M2CAI16 with the standard 10-second relaxed boundary between some of the surgical phases [54]. Also to be able to compare with previous literature [57, 35, 4], we do not use a relaxed boundary for AutoLaparo and GraSP.
For surgical tool presence detection, we use a linear head for all datasets.
For surgical action recognition we use a different head for each dataset. To evaluate on CholecT50, we utilize a linear head. The predictions are done per-frame, no temporal information is used, as in the approaches that we benchmark against [44, 43]. For SAR-RARP50, we evaluate our model in two different modes: online, where only previous frames are available (TCN head [12]), and offline ($\pm$ subscript in Table 5), which incorporates information from both previous and future frames (MS-TCN head [33]), to ensure a fair comparison with all competing methods in [48].
In our evaluation, we compare three types of models: generalist (G), pretrained on ImageNet and fine-tuned for each downstream task; specialist (S), designed for specific downstream tasks; and foundational (F), pretrained in a self-supervised manner and fine-tuned on each downstream task.
Table 4: Performance evaluation on surgical tool presence detection. We evaluate the performance on two datasets: Cholec80 [54], and GraSP [4]. The metrics are computed as described in Sec. 5.1. The evaluation protocol and the meanings of G, S, and F in the type column are described in Sec. 5.2. The superscript denotes that the backbone is frozen during fine-tuning.
Method Type Cholec80 GraSP
mAP (%) mAP (%)
ConvNeXt-L [36] G 69.3 $68.1$
ConvNeXt-L [36] $93.3$ $89.5$
EndoNet [54] S $81.0$
MTRCNet-CL [27] $89.1$
Rendezvous [44] - $80.4$
TAPIS [4] - $\mathbf{95.1}$
Ramesh et al. [49] F $93.6$
SurgFM (ours) $76.1$ $76.4$
SurgFM (ours) $\mathbf{93.7}$ $94.4$
6 Results and Discussion

### v3

@@ 3.1 Dataset Curation
Video annotation. We cross-referenced the video titles with the predefined list of procedures. When exact title matches were not found, we leveraged the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt (see Fig. 2 in supplementary material). All annotations including the procedure type and the surgery type (robotic or non-robotic) were manually reviewed and validated to ensure the accuracy of the labels assigned.
Table 2: Linear probing results for surgical phase recognition, surgical tool presence detection, and surgical action recognition tasks. The experiments are conducted with freezing the backbone and only fine-tuning a linear classification head to demonstrate the foundation model’s effectiveness. The predictions are computed on a frame-by-frame basis for all tasks.
Method Pretraining data Phase recognition Tool detection Action recognition
AutoLaparo Cholec80 M2CAI16 Cholec80 GraSP CholecT50
Acc/F1 Acc/F1 Acc/F1 mAP mAP mAP
MAE [21] LEMON 35.5/32.035.5/32.035.5 / 32.0 54.9/43.454.9/43.454.9 / 43.4 40.7/38.140.7/38.140.7 / 38.1 30.630.630.6 45.245.245.2 24.824.824.8
Endo-FM [53] Private data 51.5/43.151.5/43.151.5 / 43.1 62.7/53.962.7/53.962.7 / 53.9 53.8/49.853.8/49.853.8 / 49.8 50.050.050.0 64.364.364.3 26.226.226.2
EndoViT [5] Merged public data 45.4/38.745.4/38.745.4 / 38.7 46.9/39.046.9/39.046.9 / 39.0 38.5/37.238.5/37.238.5 / 37.2 34.734.734.7 47.147.147.1 29.529.529.5
GSViT [48] GenSurgery 22.0/17.522.0/17.522.0 / 17.5 20.2/13.520.2/13.520.2 / 13.5 22.4/18.022.4/18.022.4 / 18.0 18.218.218.2 35.635.635.6 16.516.516.5
SurgeNetXL [25] SurgeNetXL 68.8/57.068.8/57.068.8 / 57.0 73.2/65.173.2/65.173.2 / 65.1 67.8/61.967.8/61.967.8 / 61.9 72.172.172.1 62.762.762.7 45.345.345.3
LemonFM LEMON $\mathbf{76.4/66.9}$ $\mathbf{75.8/68.6}$ $\mathbf{68.4/62.4}$ $\mathbf{76.8}$ $\mathbf{76.4}$ $\mathbf{50.4}$
@@
Figure 4: Proposed augmented distillation method. To encourage invariance in LemonFM to minor surgical motion and subtle appearance changes across patients, we introduce $W_{i}$ (Eq. 1). The key component of $W_{i}$ is a pair of images (c) that are randomly selected from an augmentation pool (b). To populate the augmentation pool, which has capacity for four images, we first retrieve the nearest neighbors(1) of the input image from other videos of the same procedure type, but only include them in the pool if the cosine distance to the input image is smaller than 3×3$\times$3 × the distance between the input image and its preceding frame in the video. When not enough suitable neighbors are found, we supplement the pool with adjacent video frames(2)(3).
3.2 Ground truth and proposed new downstream tasks
Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options; and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically. The rationale for these annotations is two-fold. Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures. Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our LemonFM foundation model, video classification/retrieval, video similarity measurement, and procedure-specific estimation of surgical remaining time).
4 Proposed LemonFM Foundation Model
In this section, we describe how we build LemonFM, our foundation model pretrained on LEMON using a novel augmented distillation approach. Following the standard practice in the field of surgical computer vision [51, 52, 43], we extract frames from the 4194 videos in LEMON at 1 fps for our foundation model pretraining.
Proposed augmented distillation method. A foundation model for surgery should learn invariance to 1) minor movements of instruments and tissues observed between adjacent video frames, and 2) minimal appearance changes across patients undergoing the same surgical procedure (e.g., organ color variations). This is to encourage the model to produce similar embeddings for frames containing the same surgical tools and exhibiting comparable tissue appearances.
Inspired by self-supervised knowledge distillation [8], we introduce a novel augmented distillation approach that leverages LEMON and employs ConvNeXt-L [36] for both the teacher and student networks. Specifically, we train a student network $f_{\boldsymbol{\theta_{s}}}$ to solve for the mapping $\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))=\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ where $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ is an input image, $\sigma$ is the softmax function, $f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})\in\mathbb{R}^{C}$ is a target vector of dimension $C=2^{16}$ estimated from a teacher network $f_{\boldsymbol{\theta_{t}}}$. The student’s parameters $\boldsymbol{\theta_{s}}$ are updated by gradient descent, while the teacher’s parameters $\boldsymbol{\theta_{t}}$ are updated via exponential moving average from the student. We use the outputs of both networks to build two categorical distributions, $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ and $P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))$, where $\sigma_{z}$ denotes the $z^{th}$ element of the softmax output, and $z\in\{1,\dots,C\}$. In training, we minimize the negative log-likelihood of $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))$ following
$\min_{\boldsymbol{\theta_{s}}}\sum_{i=1}^{I}$ $\sum_{\begin{subarray}{c}\boldsymbol{u}\in U_{i}\\ \boldsymbol{v}\in V_{i}\cup W_{i}\\ \boldsymbol{u}\neq\boldsymbol{v}\end{subarray}}\Big{[}d(\boldsymbol{u},\boldsymbol{v})-n(\boldsymbol{u},\boldsymbol{v})\Big{]}$ (1)
where $d(\boldsymbol{u},\boldsymbol{v})=\log\sum_{k^{\prime}=0}^{C-1}\exp\left(\sigma_{k^{\prime}}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))\right)$, $n(\boldsymbol{u},\boldsymbol{v})=\sigma_{\arg\max_{z}P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{u})))}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))$, $I$ is the number of images in the dataset, $U_{i}$ contains two global augmentations of $\boldsymbol{x_{i}}$, and $V_{i}$ contains two global augmentations and four augmented crops of $\boldsymbol{x_{i}}$. Additionally, we introduce $W_{i}$ as an extra supervisory signal which contains two images, and four augmented crops of these two images. The procedure for selecting these two images is depicted in Fig. 4. We use the $k$-nearest neighbors (in embedding space) of the input image from other videos of the same procedure type to learn invariance to appearance changes across patients, and the adjacent video frames of the input image to learn invariance to the motion occurring between consecutive frames.
Table 3: Full fine-tuning results for surgical phase recognition, surgical tool presence detection, and surgical action recognition tasks. All experiments are conducted in a full fine-tuning setting (i.e., both backbone and head are fine-tuned). Type S denotes a specialist model trained for a specific task, whereas F denotes a foundation model whose generalization is assessed across multiple surgical downstream tasks. The superscript indicates the use of the standard 10-second relaxed boundary in phase transitions [51]. We also report the performance of LemonFM when trained on only 50% of the labeled training data for the downstream tasks.
Method Type Phase recognition Tool detection Action recognition
AutoLaparo Cholec80 M2CAI16 Cholec80 GraSP CholecT50
Acc/Jacc Acc/Jacc Acc/Jacc mAP mAP mAP
TeCNO [12] S 77.3/50.777.3/50.777.3 / 50.7 88.6/75.188.6/75.188.6 / 75.1 86.1/74.486.1/74.486.1 / 74.4 -- -- --
Trans-SVNet [30] 78.3/50.778.3/50.778.3 / 50.7 90.3/79.390.3/79.390.3 / 79.3 87.2/74.787.2/74.787.2 / 74.7 -- -- --
LoViT [35] 81.4/55.981.4/55.981.4 / 55.9 92.4/81.292.4/81.292.4 / 81.2 /-/-- / - -- -- --
MAE [21] F 63.5/42.263.5/42.263.5 / 42.2 84.6/71.284.6/71.284.6 / 71.2 77.8/63.677.8/63.677.8 / 63.6 72.372.372.3 53.053.053.0 33.133.133.1
Endo-FM [53] 80.3/55.380.3/55.380.3 / 55.3 86.9/76.786.9/76.786.9 / 76.7 79.1/69.979.1/69.979.1 / 69.9 88.488.488.4 84.284.284.2 54.754.754.7
EndoViT [5] 76.1/54.276.1/54.276.1 / 54.2 86.7/75.386.7/75.386.7 / 75.3 78.5/70.078.5/70.078.5 / 70.0 83.983.983.9 77.477.477.4 52.252.252.2
GSViT [48] 60.5/37.060.5/37.060.5 / 37.0 75.3/57.875.3/57.875.3 / 57.8 55.8/39.655.8/39.655.8 / 39.6 34.334.334.3 37.737.737.7 26.326.326.3
SurgeNetXL [25] 80.0/55.180.0/55.180.0 / 55.1 84.4/72.084.4/72.084.4 / 72.0 68.5/58.868.5/58.868.5 / 58.8 86.586.586.5 83.883.883.8 57.557.557.5
LemonFM (ours) - 50% shot 84.9/62.484.9/62.484.9 / 62.4 92.1/78.792.1/78.792.1 / 78.7 86.1/77.786.1/77.786.1 / 77.7 91.591.591.5 89.489.489.4 60.160.160.1
LemonFM (ours) $\mathbf{85.5/64.8}$ $\mathbf{92.7/85.1}$ $\mathbf{89.9/79.4}$ $\mathbf{93.7}$ $\mathbf{94.4}$ $\mathbf{61.9}$
@@ 4 Proposed LemonFM Foundation Model
Proposed video classification model. Considering that each surgical procedure is localized to a specific region of the body, we posit that a characteristic scene can be associated with each procedure. Based on this assumption, and inspired by active learning [19], we propose LemonFM-Vid. This is a video classification head that employs the frame embeddings produced by LemonFM to solve our proposed video classification downstream tasks. It does so by aggregating each frame’s embedding following
$\boldsymbol{v_{e}}=\sum_{j=0}^{J-1}\omega_{j}\boldsymbol{\phi}_{j}$ (2)
@@
5 Experiments
5.1 Downstream Tasks, Datasets and Evaluation
@@
We evaluate LemonFM on four surgical downstream tasks: surgical phase recognition [52, 51, 50], surgical tool presence detection [51, 4], surgical action recognition [43], and surgical semantic segmentation [23]. To do so, we use the most widely adopted datasets, data splits, and evaluation metrics.
Surgical phase recognition. For this task, we use three datasets: AutoLaparo [52], M2CAI16 [51], and Cholec80 [51]. These datasets include complete surgical workflows with frame-level phase annotations, where each frame corresponds to a single phase. We perform experiments with both linear probing, where we freeze the backbone and only fine-tune a linear classification head, and full fine-tuning, where both the backbone and the head are fine-tuned. We report the frame-wise accuracy and F1-score for the linear probing setting (Table 2), and the video-level accuracy and Jaccard (more details in supplementary material) for the full fine-tuning setting. To produce video-level online predictions, we replace the linear head with an MS-TCN head [12] (Table 3).
Surgical tool presence detection. Cholec80 and GraSP [4] are utilized in the surgical tool presence detection task. These datasets contain surgical tool presence labels for each frame, allowing multiple tools to be present in a single frame. We report mean Average Precision (mAP) across all frames for both the linear probing and full fine-tuning settings (see Table 2 and 3). We use a linear head for both experiments.
Surgical action recognition. We use CholecT50 [43] as the downstream dataset. Surgical action labels are annotated at frame level, enabling multiple actions occur simultaneously. We use a linear head for both the linear probing and full fine-tuning settings and apply the same metrics as in the surgical tool presence detection task.
Surgical semantic segmentation. We employ the CholecSeg8k [23] dataset, which contains pixel-wise semantic annotations for each frame. The experiments are conducted in full fine-tuning using a feature pyramid network decoder [32]. Dice score is reported over all frames in the dataset (see Table 4).
Table 4: Evaluation on surgical semantic segmentation. The experiments are conducted with using a feature pyramid network (FPN) decoder [32] in a full fine-tuning manner. The superscript denotes the LemonFM backbone is frozen. We also report the performance of LemonFM when trained on only 50% of the labeled training data.
Method Pretraining data CholecSeg8k
mDice
MAE LEMON 68.568.568.5
Endo-FM Private data 54.554.554.5
EndoViT Merged public data 73.073.073.0
GSViT GenSurgery 53.453.453.4
SurgeNetXL SurgeNetXL 70.070.070.0
LemonFM LEMON 71.971.971.9
LemonFM - 50% shot LEMON 78.278.278.2
LemonFM LEMON $\mathbf{81.3}$
5.2 Comparison with SotA
The performance of our LemonFM compared to the state of the art in each downstream task and dataset is shown in Fig. 1, and Tables 2, 3 and 4. These results demonstrate that the diversity of procedures in LEMON enables LemonFM to effectively extract features that are useful for different surgical procedures and downstream tasks.
Surgical phase recognition. As shown in Table 3, LemonFM outperformed every other surigcal foundation model in AutoLaparo ($>$9pp Jacc.), Cholec80 ($>$8pp Jacc.), and M2CAI16 ($>$9pp Jacc.). Furthermore, it also surpassed state-of-the-art specialist models for these datasets. The advantage of LemonFM persists under the linear probing setting (Table 2), showing that the pretrained features are inherently informative and transfer well without additional fine‑tuning.
Surgical tool presence detection. As shown in Table 2 and 3, LemonFM outperformed all surgical foundation models in Cholec80 and GraSP.
Surgical action recognition. LemonFM outperformed all surgical foundation models in CholecT50 (Table 2 and 3).
Surgical semantic segmentation. As shown in Table 4, LemonFM surpassed every other foundation model in the pixel-wise semantic segmentation task by a large margin ($>$8pp mDice).
Low data regime. As shown in Tables 3 and 4, when trained on 50% of the data, LemonFM outperforms other surgical foundation models fine-tuned on 100% of the data, demonstrating high data efficiency.